OpenStudy (anonymous):

A sample of gas occupies a volume of 7.50 L at 0.988 atm and 28.0 °C. (a) Calculate the pressure of the gas if its volume is decreased to 4.89 L while its temperature is held constant. (b) At what temperature in degrees Celsius is the volume of the gas 4.0 L if the pressure is kept constant.

OpenStudy (anonymous):

Use the combined gas law: \[\frac{P _{1}V _{1} }{ T _{1} } = \frac{P _{2}V _{2} }{ T _{2} } \] where: P = Pressure (in Pa) V = Volume (in m^3) T = Temperature (in K) X1 = Initial X X2 = Final X